# SnapKV [Uncaptioned image] : LLM Knows What You are Looking for Before Generation

### v1

@@ 4.2 Efficient Clustering via Pooling
In LLMs, information retrieval and generation rely on features with high attention weight and are supplemented by copying the rest in context using induction heads [15]. Hence, naively selecting the top features results in retaining only portions of details and then losing the completeness of the information. For example, such compression might cause the LLMs to retrieve only the country code of a phone number and hallucinate the rest. Our experiment also revealed that only selecting the features with the highest weights is insufficient (Sec. 5.2). Such sparse selection risks compromising the contextual integrity encapsulated in between features, thereby reducing accuracy. Based on the insights, We propose a fine-grained clustering algorithm utilizing a pooling layer shown in Line 14.
@@ 5 Experiments
In our experimental setup, we explore the performance of SnapKV across models that can handle extended sequence contexts. First, we deliver a pressure test and benchmark the speed of LWM-Text-Chat-1M [16], which is state-of-the-art regarding its context length. We then conduct an ablation study on Mistral-7B-Instruct-v0.2 to understand the influence of pooling on the model’s information retrieval performance. We assess model performances using the LongBench [17] dataset. Further, we dive into a comprehensive examination of the Command-R [2] model, another leading open-source model in the field. Lastly, we show that SnapKV can be utilized with other acceleration strategies such as parallel decoding.
@@ 5.1 Benchmarks on LWM-Text-Chat-1M
LWM-Text-Chat-1M [16] is a 7B instruction-finetuned model with up to one million context length. In this section, we conduct a pressure test on this model and examine its algorithmic efficiencies through the lens of hardware optimization.
@@ 5.1.1 Needle-in-a-Haystack
The Needle-in-a-Haystack test [18] challenges the model to accurately retrieve information from a specific sentence("needle") hidden within a lengthy document (the "haystack"), with the sentence placed at a random location. To rigorously evaluate SnapKV’s capabilities, we extended the document length to 380k tokens which is the longest content that can be processed by a single A100-80GB GPU. We configured the prompt KV cache size to 1024, enabling SnapKV to select the most crucial 1024 attention features from the prompt using our algorithm for answer generation, with a maximum pooling kernel size of 5 and a observation window size of 16. The compelling outcomes in Fig. 5 from the Needle-in-a-Haystack test underscore SnapKV’s potential to precisely manage small details on extremely long input contexts with a 380x compression ratio.
Figure 5: Needle-in-a-Haystack test performance comparison on single A100-80GB GPU, native HuggingFace implementation with only a few lines of code changed. The x-axis denotes the length of the document (the “haystack”); the y-axis indicates the position that the “needle” (a short sentence) is located within the document, from 1K to 380K tokens. For example, 50% indicates that the needle is placed in the middle of the document. Here LWMChat with SnapKV is able to retrieve the needle correctly before 160k and with only a little accuracy drop after. Meanwhile, the original implementation encounters OOM error with 33k input tokens.
Figure 6: Deconding speed comparison of baseline implementation and SnapKV optimized solutions on various batch sizes. The x-axis denotes the input sequence length; the y-axis indicates decoding speed (ms/token). All experiments are conducted on an A100 80GB GPU. The red dotted line denotes the current state-of-the-art open-sourced models’ context length.
@@ 5.1.2 Decoding Speed and Memory Bound
We further benchmark the speed of LWM-Text-Chat-1M under different batch-size settings using SnapKV. We set the maximum prompt KV cache size as 2048 for SnapKV. There are two main takeaways from our experiment on decoding speed and input sequence length on various batch sizes, as shown in Fig. 6. First, as the input sequence length increases, the decoding speed of the baseline implementation escalates exponentially. Conversely, the SnapKV-optimized model maintains a constant decoding speed since the KV cache stays the same and there is no extra update during the inference. For instance, at a sequence length of 16k and a batch size of 2, the decoding time for the baseline model surpasses 0.1 seconds, whereas the SnapKV-optimized model consistently remains below 0.04 seconds, achieving approximately a 3.6x speedup. Second, with the same batch size, the model optimized with SnapKV can decode significantly longer sequences. For example, at a batch size of 2, the baseline model encounters an OOM issue beyond 16k input tokens, whereas the SnapKV-enhanced model extends this limit to 131k input tokens, indicating an approximately 8.2x improvement. This demonstrates SnapKV’s effectiveness in minimizing memory consumption.
@@ 5.2 Ablation Study of Effectiveness of Pooling
Refer to captionRefer to caption
We perform an ablation study to assess the impact of our pooling technique, a straightforward but efficient method for consolidating information through clustering. Our evaluation utilizes the modified LongEval-Lines benchmark [19], incorporating random generated pairs and averaged scores. LongEval-Lines presents a greater challenge compared to Needle-in-a-Haystack because it involves identifying key-value pairs in noisy contexts of the same format, while in Needle-in-a-Haystack, the relevant information is more distinctly separated from other contexts. We apply max pooling with a kernel size of 5 and use the observation window with a size of 16. The findings, illustrated in our results (Fig. 7), indicate that pooling significantly enhances retrieval accuracy compared to methods not utilizing pooling. We hypothesize that this is due to the ability of strong attention mechanisms to focus on the initial portion of tokens. Without information compression, large language models tend to replicate the subsequent tokens, leading to retrieved partially correct results when the KV cache is compressed as we observed. Note that throughout our experiments, the choice between max pooling and average pooling did not yield significant differences in performance.
@@
   LLMsa Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code
@@
  • a Credit to Jin et al. [20] for the template used in the table.
@@ 5.3 Experiments on LongBench
We evaluate SnapKV on these four models using LongBench [17], a multi-task benchmark designed to rigorously evaluate long context understanding capabilities across various datasets, spanning single and multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion. We choose LWM-Text-Chat-1M with 1 million context length, LongChat-7b-v1.5-32k, Mistral-7B-Instruct-v0.2, Mixtral-8x7B-Instruct-v0.1 with 32k context length as our baselines. For each model, we test SnapKV with various settings: compressing KV caches in the prompt to 1024, 2048, and 4096 tokens. We use max pooling with kernel size 7 and observation window size 32. Table 1 illustrates a negligible performance drop from models with SnapKV compared with original implementations for 16 different datasets, even with prompt-KV with 1024 tokens. Some models even outperform the baseline. Our results substantiate that SnapKV can grasp the key information in the long context and give comprehensive summaries with details. Moreover, our results also indicate the effectiveness of SnapKV in compressing the prompt KV cache. For LongChat-7b-v1.5-32k, the average input token length is 12521; for LWM-Text-Chat-1M, 13422; for Mistral, 13160. Thus, using 1024, SnapKV achieves an average compression rate of 92%, and using 4096, it reaches 68%, all with negligible drops in accuracy. We compare SnapKV and H2O on the LongBench dataset to further demonstrate the performance of SnapKV. To fairly evaluate the accuracy, we set the prompt capacity for H2O to 4096. As table 1 shows, SnapKV delivers significantly better performance than H2O. Even with 1024 prompt KV caches, SnapKV on Mistral-7B-Instruct-v0.2 achieves better performance than H2O with 4096 caches on 11 out of 16 benchmarks.
@@ 6 Discussions
SnapKV emerges as a potent yet straightforward solution, adeptly compressing the KV caches of models to mitigate the computational and memory burdens associated with processing extensive inputs. Originating from a nuanced observation that specific tokens within prompts garner consistent attention from each head during generation, our methodology not only conserves crucial information but also enhances processing efficiency. Despite its strengths, SnapKV’s scope is primarily confined to the generative aspect of models, specifically targeting the KV caches during the generation. This limitation implies that SnapKV cannot extend a model’s long context capability if the model inherently struggles with long contexts or exhibits poor performance. Additionally, SnapKV’s design does not cover the processing of the prompt inference, which limits its effectiveness in scenarios where the system cannot handle prompts of extensive length. Nonetheless, our contributions offer significant insights and tools for the community, paving the way for more refined approaches on managing the challenges of large-scale language modeling.
Appendix A Visulization of the Generated Context
[Uncaptioned image][Uncaptioned image]

### v2

@@ 4.3 Efficient Clustering via Pooling
In LLMs, information retrieval and generation rely on features with high attention weight and are supplemented by copying the rest of features in context using induction heads [15]. Hence, naively selecting the top features results in retaining only portions of details and then losing the completeness of the information. For example, such compression might cause the LLMs to retrieve only the country code of a phone number and hallucinate the rest. Our experiment also revealed that only selecting the features with the highest weights is insufficient (Sec. 5.2). Such sparse selection risks compromising the contextual integrity encapsulated in between features, thereby reducing accuracy. Based on the insights, we propose a fine-grained clustering algorithm utilizing a pooling layer shown in Line 13.
@@ 5 Experiments
In our experimental setup, we explore the performance of SnapKV across models that can handle extended prompt sequence contexts. First, we deliver a pressure test and benchmark the speed of LWM-Text-Chat-1M [16], which is state-of-the-art regarding its context length. We then conduct an ablation study on Mistral-7B-Instruct-v0.2 to understand the influence of pooling on the model’s information retrieval performance. We assess model performances using the LongBench [17] dataset. Further, we dive into a comprehensive examination of the Command-R [2] model, another leading open-source model in the field. Lastly, we show that SnapKV can be utilized with other acceleration strategies such as parallel decoding.
@@ 5.1 Benchmarks on LWM-Text-Chat-1M
LWM-Text-Chat-1M [16] is a 7B instruction-fine-tuned model with up to one million context length. In this section, we conduct a pressure test on this model and examine its algorithmic efficiencies.
Figure 6: Needle-in-a-Haystack test performance comparison on single A100-80GB GPU, native HuggingFace implementation with only a few lines of code changed. The x-axis denotes the length of the document (the “haystack”) from 1K to 380K tokens; the y-axis indicates the position that the “needle” (a short sentence) is located within the document. For example, 50% indicates that the needle is placed in the middle of the document. Here LWMChat with SnapKV is able to retrieve the needle correctly before 140k and with only a little accuracy drop after. Meanwhile, the original implementation encounters OOM error with 33k input tokens (white dashed line).
@@ 5.1.1 Needle-in-a-Haystack
The Needle-in-a-Haystack test [18] challenges the model to accurately retrieve information from a specific sentence ("needle") concealed within an extensive document (the "haystack"), with the sentence placed at a random location. Typically, sentences that are inserted in the middle of prompts are harder to retrieve. To rigorously evaluate SnapKV’s capabilities, we extended the document length to 380k tokens which is the longest content that can be processed by a single A100-80GB GPU. We configured the prompt KV cache size to 1024, enabling SnapKV to select the most crucial 1024 attention features from the prompt for answer generation, with a maximum pooling kernel size of 5 and an observation window size of 16, both of which are hyperparameters that can be customized. The compelling outcomes in Fig. 6 from the Needle-in-a-Haystack test underscore SnapKV’s potential to precisely manage small details on extremely long input contexts with a 380x compression ratio.
Figure 7: Decoding latency comparison of baseline implementation and SnapKV optimized solutions on various batch sizes. The x-axis denotes the input sequence length; the y-axis indicates decoding latency (ms/token). All experiments are conducted on an A100 80GB GPU. The red dotted line denotes the common context length of state-of-the-art long sequence models.
@@ 5.1.2 Decoding Speed and Memory Bound
We further benchmark the speed of LWM-Text-Chat-1M under different batch-size settings using SnapKV. We set the maximum KV cache size as 2048 for SnapKV, and fix the generation length at 512 to ensure a fair comparison. There are two main takeaways from our experiment on decoding speed and prompt sequence length on various batch sizes, as shown in Fig. 7. First, as the input sequence length increases, the decoding latency of the baseline implementation escalates linearly. Conversely, the SnapKV-optimized model maintains a constant decoding speed since the compressed KV cache size of prompt stays the same regardless of input sequence length and there is no extra update during the inference. For instance, at a sequence length of 16k and a batch size of 2, the decoding time for the baseline model surpasses 100 ms, whereas for SnapKV-optimized model, the decoding time consistently remains below 40 ms, achieving approximately a 3.6x speedup. Second, with the same batch size, the model integrated with SnapKV can decode significantly longer sequences. For example, at a batch size of 2, the baseline model encounters an OOM error beyond 16k input tokens, whereas the SnapKV-enhanced model extends this limit to 131k input tokens, indicating an approximately 8.2x improvement. This demonstrates SnapKV’s effectiveness in minimizing memory consumption.
@@ 5.2 Ablation Study of Effectiveness of Pooling
We perform an ablation study on Mistral-7B-Instruct-v0.2 to assess the impact of our pooling technique, a straightforward but efficient method for consolidating information through clustering. Our evaluation utilizes the modified LongEval-Lines benchmark [19], incorporating randomly generated pairs and averaged scores. LongEval-Lines presents a greater challenge compared to Needle-in-a-Haystack because it involves identifying key-value pairs in noisy contexts of the same format, while in Needle-in-a-Haystack, the relevant information is more distinctly separated from other contexts. We apply max pooling with a kernel size of 5 and use the observation window with a size of 16, which are hyperparameters and could be customized according to different models. As illustrated in our results (Fig. 8), we find that pooling significantly enhances retrieval accuracy compared to methods not utilizing pooling. We hypothesize that this is because the initial portions of critical token clusters are weighted higher by attention mechanisms. Typically, large language models tend to copy the tokens surrounding the initial portions to keep the contextual integrity. However, naively compressed KV cache breaks this mechanism and could lead to partially correct results (Fig. 8). Note that throughout our experiments, the choice between max pooling and average pooling did not yield significant differences in performance.
@@
   LLMs * Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code
@@
  • * Credit to Jin et al. [20] for the template used in the table.
@@ 5.3 Experiments on LongBench
We evaluate SnapKV on these four models using LongBench [17], a multi-task benchmark designed to rigorously evaluate long context understanding capabilities across various datasets, spanning single and multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion. We choose LWM-Text-Chat-1M with 1 million context length, LongChat-7b-v1.5-32k, Mistral-7B-Instruct-v0.2, Mixtral-8x7B-Instruct-v0.1 with 32k context length as our baselines. For each model, we test SnapKV with various settings: compressing KV caches in the prompt to 1024, 2048, and 4096 tokens. We use max pooling with kernel size 7 and observation window size 32. Table 1 illustrates a negligible performance drop from models with SnapKV compared with original implementations for 16 different datasets, even with prompt-KV with 1024 tokens. Some models even outperform the baseline. Our results substantiate that SnapKV can grasp the key information in the long context and give comprehensive summaries with details. Moreover, our results also indicate the effectiveness of SnapKV in compressing the prompt KV cache. For these 4 models, the average input token length is around 13k. Thus, using 1024, SnapKV achieves an average compression rate of 92%, and using 4096, it reaches 68%, all with negligible drops in accuracy. We compare SnapKV and H2O on the LongBench dataset to further demonstrate the performance of SnapKV. To fairly evaluate the accuracy, we set the prompt capacity for H2O to 4096. As Table 1 shows, SnapKV delivers significantly better performance than H2O. Even with 1024 prompt KV caches, SnapKV on Mistral-7B-Instruct-v0.2 achieves better performance than H2O with 4096 caches on 11 out of 16 benchmarks.
@@ 6 Discussions
SnapKV is an effective yet straightforward solution that compresses the KV cache to mitigate the computational and memory burdens of processing extensive prompts. Observing that specific tokens within prompts gain consistent attention from each head during generation, our methodology not only retrieve crucial information but also enhances processing efficiency. Despite its strengths, SnapKV’s scope is primarily confined to the generative aspect of models, specifically targeting the KV cache during the generation. This limitation implies that SnapKV cannot extend a model’s long context capability if the model inherently struggles with long contexts or exhibits poor performance. Additionally, SnapKV’s design does not cover the processing of the prompt inference, which limits its effectiveness in scenarios where the system cannot handle prompts of extensive length. Nonetheless, our contributions offer significant insights and tools for the community, paving the way for more refined approaches on managing the challenges of large-scale language modeling. The appendix provides more experiments with parallel decoding and the discussion about generation speedup.
Appendix A Discussion of Generation Time Speedup
To better assess SnapKV’s effectiveness across different stages, we documented a detailed time breakdown for Mistral-7B-Instruct-v0.2 during both the prompting and generation stages. We configured the model to consistently generate 512 tokens, facilitating a direct comparison with the prompting stage. As illustrated in Figure 10, generation time dominates the whole processing time for LLMs over input sequences, introducing significant overhead. While the generation time for the original model increases with input length, SnapKV maintains a consistent decoding speed regardless of input length, significantly reducing generation time. Especially, SnapKV is able to achieve balanced prompting time and generation time with input length smaller than 100k.
Figure 10: The prompting time and generation time comparison between Mistral model with and without SnapKV.
Appendix B Visulization of the Generated Context